OpenStudy (anonymous):

Find the percent of increase from 320 to 380. Round to the nearest tenth of a percent if necessary. (2 points)

OpenStudy (anonymous):

320 * x = 380 x = 380/320 380/320 - 1 = 0.1875 so it's 18.8%